OpenStudy (anonymous):

1. Simplify 3fg(f2g + 5f – 4). 3f2g2 +15fg – 12 3f3g3 + 15f2g – 12 3f3g2 + 15f2g – 12 3f3g2 + 15f2g – 12fg Tell whether each sequence is arithmetic, geometric, or neither. Find the next three terms of each sequence. Write a rule to describe the sequence. 2. –3, –12, –48, ... –192; –768; –3,072; geometric; Start with –3 and multiply by 4 repeatedly. –198; –788; –3,072; arithmetic; Start with 4 and multiply by –3 repeatedly. –188; –768; –3,058; arithmetic; Start with 4 and multiply by –3 repeatedly. –194; –788; –3,072; geometric; Start with 3 and multiply by 4 repeated

OpenStudy (anonymous):

\[3f^3g^2+15f^2g-12fg\] so the last one

OpenStudy (anonymous):

now the sequence...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you for helping me:D

OpenStudy (anonymous):

sorry, called into a meeting will get back to it in about 1 hr, sorry

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

back for a sec :D Question 2 is the top answer. -192,-168, -3072 and you tell its geometric because the ratio of the a and a+1 is 0.25 i.e. -3/-12 = 0.25 -12/-48 = 0.25 as well

OpenStudy (anonymous):

good luck :D

OpenStudy (anonymous):

the last thing it asks for is the rule: this follows the geometric sequence rule a(n)=a(1)*r^(n-1) so we have a(n) = -3*4^(n-1) is four because the the terms have a fixed ration of (a+1)/a or -12/-3 = 4